Three times the complement of an angle is twenty more than the supplement. Find the angles, the complement, and the supplement.
Answer by ankor@dixie-net.com(22740)

Three times the complement of an angle is twenty more than the supplement.
Find the angles, the complement, and the supplement.
:
Let A = "the angle"
the equation for the statement:
"Three times the complement of an angle is twenty more than the supplement."
3(90-A) = (180-A) + 20
270 - 3a = 200 - A
270 - 200 = -A + 3A
70 = 2A
A = 70/2
A = 35 degrees is the angle
then
90 - 35 = 55 degrees is it's complement
and
180 - 35 = 145 degrees is it's supplement
:
:
Check this in the statement
"Three times the complement of an angle is twenty more than the supplement."
3(55) = 145 + 20
165 = 165